# Gauge Symmetries, Contact Reduction, and Singular Field Theories

### v1

@@ Abstract
The reduction of dynamical systems which are invariant under changes of global scale is well-understood, for classical theories of particles, and fields. The excision of the superfluous degree of freedom describing such a scale leads to a dynamically-equivalent theory, which is frictional in nature. In this article, we extend the formalism to physical models, of both particles and fields, described by singular Lagrangians. Our treatment of classical field theory is based on the manifestly covariant Hamilton De-Donder Weyl formalism, in which the Lagrangian density is introduced as a bundle morphism on the pre-multisymplectic velocity phase space $J^{1}E$. The results obtained are subsequently applied to a number of physically-motivated examples, as well as a discussion presented on the implications of our work for classical General Relativity.
@@ I Introduction
Singular field theories are, in many cases, amongst the most interesting class of models encountered in the description of natural phenomena [1, 2, 3]. Indeed, the most precisely tested mathematical theory to date, the Standard Model of Particle Physics, is a chiral quantum field theory, gauged under local transformations of the group $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$ [4, 5, 6]. Further, it is postulated that any UV-complete theory of quantum gravity may possess no global symmetries, and that all symmetries must therefore be either spontaneously broken, or gauged [7, 8, 9]. It is thus clear that singular theories are of central importance to the advancement of modern theoretical physics.
Any attempt to rigorously develop a quantum theory requires a firm understanding of the classical counterpart upon which it is based. As such, in this article, we shall present a framework in which singular theories that possess scaling symmetries may be reduced to a lower-dimensional, dynamically-equivalent description, that is frictional in nature [10]. For theories of particles, we take, as our starting point, the geometrical constraint algorithm for pre-symplectic systems, developed by Nester et. al. in [11, 12, 13]. This is presented in section (III.1), following a brief review of elementary pre-symplectic geometry in section (II).
@@ II Geometrical Preliminaries
Suppose that $Q$ is an $n$-dimensional smooth manifold, corresponding to the configuration space of some mechanical system. Local coordinates on $Q$ are taken to be $q^{i}$, and we introduce a Lagrangian function $L:TQ\rightarrow\mathbb{R}$ on the tangent bundle [35, 36, 37]. Fibre coordinates of $TQ$ are denoted $v^{i}$, and we have the bundle projection $\pi:TQ\rightarrow Q$. The tangent bundle is endowed with a number of canonical geometrical features, including the vertical endomorphism $J:T(TQ)\rightarrow T(TQ)$, satisfying $J^{2}=0$, and $\textrm{rank}\,J=n$, which constitutes an almost-tangent structure on $TQ$ [38, 39]. Additionally, we have a distinguished Liouville vector field $\Delta\in\mathfrak{X}^{\infty}(TQ)$, defined such that its one-parameter family of transformations generates dilatations of the fibres of $TQ$. In local bundle coordinates $(q^{i},v^{i})$, we have
$J=\frac{\partial}{\partial v^{i}}\otimes dq^{i}\quad\quad\quad\quad\Delta=v^{i}\,\frac{\partial}{\partial v^{i}}$ (1)
Given a Lagrangian function $L$, the canonical structures $J$ and $\Delta$ are used to define a (pre-)symplectic Lagrangian system. In particular, the function $E_{L}:=\Delta(L)-L$ corresponds to the scalar energy of the mechanical system described by $L$, whilst $\theta_{L}:=\,^{t}\hskip-1.70717ptJ\circ dL$ is used to construct, the Poincaré-Cartan form $\omega_{L}:=-\,d\theta_{L}$. Clearly, $\omega_{L}$ is closed; non-degeneracy, on the other hand, is not assured, and throughout subsequent sections, we shall be interested in those systems for which $\omega_{L}$ is degenerate. In such cases, we refer to the triple $(TQ,\omega_{L},E_{L})$ as a pre-symplectic Lagrangian system [40], and from (1) it follows that
$E_{L}=\frac{\partial L}{\partial v^{i}}v^{i}-L\quad\quad\quad\quad\omega_{L}=-\,d\left(\frac{\partial L}{\partial v^{i}}dq^{i}\right)$ (2)
@@
be of non-maximal rank, and thus singular. The equations of motion for a Lagrangian system $(TQ,\omega_{L},E_{L})$ are formulated geometrically via the following vector bundle morphism
$\begin{split}\flat_{\scriptscriptstyle L}\;:\;T(TQ)\,&\longrightarrow\,T^{*}(TQ)\\ X\;&\longmapsto\;\iota_{X}\omega_{L}\end{split}$ (3)
It should be noted that when $\omega_{L}$ is degenerate, $\flat_{\scriptscriptstyle L}$ is not an isomorphism. The dynamics are deduced seeking a vector field $X_{L}\in\mathfrak{X}^{\infty}(TQ)$, such that
$\flat_{\scriptscriptstyle L}(X_{L})=\iota_{X_{L}}\omega_{L}=dE_{L}$ (4)
When $(TQ,\omega_{L},E_{L})$ is a symplectic system, such vector fields are guaranteed to exist, and are unique. In addition, integral curves of $X_{L}$ project down to curves on the configuration space $Q$, and these latter curves correspond to physical trajectories of the mechanical system. Such vector fields are interchangeably referred to as holonomic or second-order differential equations (SODE) [41].
When the system of interest possesses a singular Lagrangian, not only are vector field solutions of (4) not guaranteed to exist, but since $\textrm{ker}\,\omega_{L}$ is non-trivial, any solution that is found is no longer unique [40]. In the most favourable of cases, there will exist a submanifold $P_{f}\hookrightarrow TQ$, upon which (4) admits solutions, and this space is deduced through use of a geometrical constraint algorithm [12, 13].
For the purposes of studying the contact reduction of singular systems, we shall work almost exclusively within the Hamiltonian formalism. Our motivation for this is twofold: on the one hand, the Hamiltonian framework provides a well-defined bracket structure, which readily allows us to classify the constraint functions, and deduce the dynamical evolution of any phase space variable. Further, not only does the Lagrangian formalism lack such a bracket, but the geometrical constraint algorithm introduces a myriad of additional complications, which arise owing to a need to impose that, at each step, the vector field solution satisfy the second-order problem. Since the dynamical content of the two formalisms is identical, we shall evade the undesirable features of the Lagrangian description by using it only as a means to obtain the corresponding Hamiltonian function.
@@
In order to formulate the dynamical problem, the vector bundle morphism $\flat:T(T^{*}Q)\rightarrow T^{*}(T^{*}Q)$ is restricted to $M_{0}$, and the resulting map denoted $\flat_{0}$. We then seek a vector field $X_{H}$, such that
@@
As in the Lagrangian case, solutions generally do not exist on the entirety of $M_{0}$, thereby necessitating the implementation of a constraint algorithm, which is described in detail in section (III.1). The final constraint manifold $M_{f}$ is then the maximal subspace of the cotangent bundle upon which the dynamical problem (5) possesses well-defined (albeit non-unique) solutions.
@@ III.1 The Hamiltonian Algorithm
When analysing the dynamics of singular systems, we shall use the almost-regular Lagrangian $L:TQ\rightarrow\mathbb{R}$ to obtain the pre-symplectic Hamiltonian system $(M_{0},\omega_{0},H_{0})$, introducing the embedding $\jmath_{0}:M_{0}\hookrightarrow T^{*}Q$. It will often be convenient to describe $M_{0}$ as the zero-set of a collection of primary constraint functions $\{\phi^{a}\}$, with $a=1,\,\cdots,n-k$, in which $n=\textrm{dim}\,Q$, and $k=\textrm{rank}\,W_{ij}$ [43]. We shall assume throughout that all subspaces generated are regular, closed submanifolds of $T^{*}Q$, whose natural injections are embeddings. We therefore begin by seeking a vector field $X_{H}$ which satisfies
@@
Since $\omega_{0}$ is pre-symplectic, $\flat_{0}$ is not an isomorphism; consequently, an initial prerequisite for our geometrical equation to possess solutions is that, at every point, $dH_{0}$ must lie in the range of the map $\flat_{0}$. Given that such a condition is not guaranteed to be satisfied a priori, the first step of the constraint algorithm requires us to restrict $M_{0}$ to the following subset
$M_{1}:=\{x\in M_{0}\;|\;(dH_{0})_{x}\in\,\flat_{0}(T_{x}M_{0})\,\}$ (9)
From the discussion above, under the assumptions of reflexivity and topological closedness, we recall that the image $\flat_{0}(T_{x}M_{0})$ coincides with the annihilator of $T_{x}M_{0}^{\perp}$; this is particularly noteworthy, as $T_{x}M_{0}^{\perp}$ is precisely $\textrm{ker}\,(\omega_{0})_{x}$, which is readily deduced in local coordinates. Thus, from a pragmatic standpoint, it is preferable to characterise $M_{1}$ as follows
@@
We refer to these objects as the secondary constraint functions. In order for any vector field solution $X_{H}$, defined over the points of $M_{1}$, to be physically meaningful, it must remain tangent to $M_{1}$. If this were not the case, the solution would tend to evolve off the constraint surface, and thus cease to obey the restrictions placed upon the system. Tangency of the solution is not guaranteed, and so must be imposed as an additional constraint, leading us to consider only those points of $M_{1}$ at which the solution may be written as $X_{H}=\jmath_{1*}Z$ for some $Z\in TM_{1}$; generally, $\jmath_{k}:M_{k}\hookrightarrow M_{k-1}$ is used to refer to the inclusion of $M_{k}$. The condition $X_{H}=\jmath_{1*}Z$ is not readily imposed locally, and so we shall use the following alternative criterion to characterise $M_{2}$
@@
That this provides an equivalent definition to that given by imposing $X_{H}=\jmath_{1*}Z$ follows from a series of simple calculations, outlined in [11]. Vector field solutions on $M_{2}$ are now tangent to $M_{1}$, but, in general, not to $M_{2}$, requiring this condition to be imposed as a further restriction. It is then clear from (9) and (12) how the algorithm must proceed: at the $k^{\textrm{th}}$ step, the submanifold $M_{k}$ is composed of all points of $M_{k-1}$ at which the solution is tangent to $M_{k-1}$, but not necessarily to $M_{k}$, leading us to restrict $M_{k}$ to the smaller set $M_{k+1}$. This condition may be imposed requiring that $X_{H}=\jmath_{k*}Z$ for some $Z\in TM_{k}$, or equivalently, by demanding that $dH_{0}$ belong to the annihilator $(TM_{k}^{\perp})^{o}$, so that
@@
The constraint procedure may terminate in one of three ways. Firstly, there may be some $N\in\mathbb{N}$, for which $M_{N}=\emptyset$; in this case, the equation of motion (8) possesses no solution and the system $(M_{0},\omega_{0},H_{0})$ does not describe any physical configuration. Alternatively, the algorithm may yield a submanifold $M_{N}$, which is non-empty, but of dimension zero; such a space may be identified as a collection of isolated points, which forces the vector field solution to be identically zero. Evidently, neither of these eventualities is particularly interesting from a physical perspective. Consequently, we shall only consider those systems for which there exists some $N\in\mathbb{N}$, such that $M_{N+1}=M_{N}$, with $\textrm{dim}\,M_{N}\neq 0$. Upon deducing the value of $N$ for which the algorithm terminates, we shall declare $M_{f}:=M_{N}$ to be the final constraint manifold.
@@ III.2 Geometry of the Final Constraint Submanifold
The algorithm of the previous section provides a systematic way to restrict the set of points of the cotangent bundle at which the equations of motion of a singular system admit everywhere-tangent solutions. At no point was it necessary to specify the nature of the constraints defining each subspace; once the final submanifold has been found, however, the dynamical evolution of observable quantities is deduced locally, which does require the distinction between first and second class constraints to be made.
Suppose that the constraint surface $M_{f}$ is defined as the zero-set of the functions $\phi^{\alpha}$, with $\alpha=1,\,\cdots,D$. We introduce the matrix $J$, whose entries are the Poisson brackets of the constraints i.e $J^{\alpha\beta}:=\{\phi^{\alpha},\phi^{\beta}\}$. If $\textrm{rank}\,J=D-F$, it follows that there exist $F$ $C^{\infty}$-linear combinations of the $\phi^{\alpha}$ which generate unobservable (gauge) transformations. [^1] Here, and throughout all subsequent analyses, we take the Dirac conjecture - that all first class constraints generate gauge transformations - to be true [44, 45, 46]. The goal is then to identify such combinations, corresponding to first class functions, and to modify the original set of constraints, so as to extract a maximal subfamily of second class functions. Note that, since $J$ is calculated from the Poisson bracket, it follows that $\textrm{rank}\,\omega_{f}=D-F$, in which $\omega_{f}:=\iota^{*}\omega$ is the closed 2-form naturally inherited by $M_{f}$, and $\imath:=\jmath_{f}\circ\cdots\circ\jmath_{0}$ defines the embedding of $M_{f}$ into $T^{*}Q$.
In order to extract the first class functions, we introduce a ‘vector’ $v_{\alpha}=(v_{1},\,\cdots,v_{D})$, and demand that
@@
It is sufficient to seek these null vectors on the constraint manifold, since a function $f$ is only required to have weakly-vanishing Poisson bracket with the $\phi^{\alpha}$, to be characterised as first class. Having identified all linearly independent combinations satisfying (14), indexing the resulting vectors via $v^{(a)}$, the $a^{\textrm{th}}$ first class constraint is simply $\Omega^{a}:=v^{(a)}_{\alpha}\phi^{\alpha}$. In order to extract the maximal subfamily of second class constraints, we seek DFD-F linearly independent vectors $w^{(j)}$, which are not in the kernel of $J$, so that
@@
The Dirac bracket has a number of noteworthy properties:
  • $\{f,gh\}_{D}=\{f,g\}_{D}\,h+g\{f,h\}_{D}$ for all $f,g,h\in C^{\infty}(T^{*}Q)$
  • $\{f,\chi^{j}\}_{D}=0$ for any $f\in C^{\infty}(T^{*}Q)$
  • $\{f,g\}_{D}\approx\{f,g\}$ for $f$ first class and $g$ arbitrary
We now impose the second class constraints $\chi^{j}=0$ as strong equalities, which defines a hypersurface
@@
with inclusion $\jmath_{\chi}:M_{\chi}\hookrightarrow T^{*}Q$, which inherits a closed 2-form $\omega_{\chi}:=\jmath^{*}_{\chi}\omega$ which is symplectic. The Dirac bracket is precisely the algebraic structure required to ‘transfer’ this symplectic structure to functions defined on the whole of $T^{*}Q$. In particular, the dynamical evolution of any phase space function $f\in C^{\infty}(T^{*}Q)$ may be deduced from $\dot{f}=\{f,H_{T}\}_{D}$, in which $H_{T}:=H+\lambda_{a}\Omega^{a}$ is the total Hamiltonian; $\lambda^{a}$ are arbitrary Lagrange multipliers, and $H$ is any extension of $H_{0}$ to the full phase space.
@@
$M_{\scriptscriptstyle\Omega}$ is an embedded submanifold of $M_{\chi}$, with inclusion $\jmath_{\scriptscriptstyle\Omega}:M_{\scriptscriptstyle\Omega}\hookrightarrow M_{\chi}$; the 2-form $\omega_{\scriptscriptstyle\Omega}:=\jmath^{*}_{\scriptscriptstyle\Omega}\,\omega_{\chi}$ is maximally degenerate, since $M_{\scriptscriptstyle\Omega}$ is first class. It is straightforward to show that the Hamiltonian vector fields $X_{a}$ corresponding to $\Omega^{a}$ via $\iota_{X_{a}}\omega_{\chi}=d\Omega^{a}$ are tangent to $M_{\scriptscriptstyle\Omega}$, since
$\mathcal{L}_{X_{a}}\Omega^{b}=d\Omega^{b}[X_{a}]=\omega_{\chi}(X_{b},X_{a})=\{\Omega^{b},\Omega^{a}\}\approx 0$
Here, the weak equality is taken to signify ‘when restricted to $M_{\scriptscriptstyle\Omega}$.’ In general, a vector field $\xi\in\mathfrak{X}^{\infty}(T^{*}Q)$ may only be tangent to $M_{\scriptscriptstyle\Omega}$ if $\mathcal{L}_{\xi}\Omega^{a}\approx 0$ for all $\Omega^{a}$: that is if $\xi\in\textrm{ker}\,\omega_{\scriptscriptstyle\Omega}$; since the first class constraints are assumed to be independent (by which we mean that their differentials are linearly independent) it follows that $\textrm{ker}\,\omega_{\scriptscriptstyle\Omega}$ is spanned by the Hamiltonian vector fields $X_{a}$.
More formally, at each point $x\in M_{\scriptscriptstyle\Omega}$, the symplectic orthogonal
$T_{x}M_{\scriptscriptstyle\Omega}^{\perp}:=\{v\in T_{x}M_{\chi}\;|\;\omega_{\chi}(v,w)=0\quad\textrm{for all}\;w\in T_{x}M_{\scriptscriptstyle\Omega}\,\}$
satisfies $T_{x}M_{\scriptscriptstyle\Omega}^{\perp}\subseteq T_{x}M_{\scriptscriptstyle\Omega}$, and so $M_{\scriptscriptstyle\Omega}$ constitutes a coisotropic submanifold of $M_{\chi}$ [48]; for such a class of space, the distribution $x\mapsto T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$ is involutive, since, for any two $X,Y\in T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$, we take some $Z\in T_{x}M_{\scriptscriptstyle\Omega}$, and compute
$d\omega_{\chi}(X,Y,Z)=X\omega_{\chi}(Y,Z)-Y\omega_{\chi}(Z,X)+Z\omega_{\chi}(X,Y)-\omega_{\chi}([X,Y],Z)+\omega_{\chi}([X,Z],Y)-\omega_{\chi}([Y,Z],X)$
Since $\omega_{\chi}$ is closed, $d\omega_{\chi}(X,Y,Z)=0$; moreover, $X$ and $Y$ were taken to be in the symplectic orthogonal $T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$, so that the first three terms on the right-hand side vanish. Similar considerations lead us to conclude that
$d\omega_{\chi}(X,Y,Z)=-\,\omega_{\chi}([X,Y],Z)=0$
so that if $X,Y\in T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$, then $X,Y]\in T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$. Hence, $x\mapsto T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$ is an involutive distribution, which, locally, we may write as the span of the $X_{a}$ [49]: that is to say
@@
Since the Hamiltonian vector fields of the first class constraints define an involutive distribution, we know, from the Frobenius theorem, that there exists a foliation $\mathcal{F}$ of $M_{\scriptscriptstyle\Omega}$ into $F$-dimensional submanifolds, which, in this context, we refer to as gauge leaves [50, 51]. The gauge orbits correspond to integral curves of the $X_{a}$, and they map points of $M_{\scriptscriptstyle\Omega}$ into other points, which produce dynamically indistinguishable configurations. Note that these results are valid only on $M_{\scriptscriptstyle\Omega}$: elsewhere, the vector fields $X_{a}$ do not, in general, close under the Lie bracket.
As a result of the degeneracy of $\omega_{\scriptscriptstyle\Omega}$, the constraint surface does not inherit a well-defined bracket structure. It is therefore necessary to construct a reduced space, quotienting out by the gauge orbits; this reduced phase space identifies all points related by gauge transformations, and thus contains only configurations which are physically inequivalent. Upon constructing this space, there arise two distinct possibilities: in the most favourable of cases, the algebra formed by the first class constraints is a genuine Lie algebra; this occurs precisely when the structure functions $f^{ab}_{c}$, appearing in $\{\Omega^{a},\Omega^{b}\}=f^{ab}_{c}\Omega^{c}$, are constant over the whole of $M_{\scriptscriptstyle\Omega}$. In these conditions, the finite gauge transformations are generated by the elements of the corresponding Lie group $G$ that are path-connected to the identity, [^2] This lifting of the elements of the Lie algebra to the group is, of course, provided by the exponential map. and they give rise to a smooth action of $G$ on $M_{\scriptscriptstyle\Omega}$. It is then possible to perform a reduction procedure, using results similar to those of the Marsden-Weinstein theorem [52, 53], to obtain a phase space $M_{\scriptscriptstyle\Omega}/G$, which is symplectic [54, 55].
In most cases, such conditions will not be realised; it is often found that the Dirac bracket between two first class constraints is strongly equal to some combination of the phase space variables which vanishes on $M_{\scriptscriptstyle\Omega}$. The structure functions are then of the form $f^{ab}_{c}(q^{i},p_{i})$, and they define a Lie algebroid on $M_{\scriptscriptstyle\Omega}$ [56]. Formally, it is again possible to quotient out by the gauge orbits, obtaining a physical phase space that is symplectic; however it is almost invariably more straightforward to introduce a choice of gauge-fixing. Since we are only concerned with the dynamical evolution of gauge-invariant observable quantities, the precise choice of gauge-fixing is somewhat arbitrary, provided that the gauge is obtainable via a finite number of infinitesimal transformations, and that it fixes all degrees of freedom, leaving no remnant gauge symmetry [57, 58].
Upon carrying out the pre-symplectic reduction process, we obtain a physical state space $\mathcal{P}$, together with a projection map $\pi:M_{\scriptscriptstyle\Omega}\rightarrow\mathcal{P}$, sending each point on the constraint surface to its equivalence class under gauge orbits. The reduced phase space $\mathcal{P}$ is symplectic, with associated 2-form $\pi^{*}\omega_{\scriptscriptstyle\mathcal{P}}=\jmath^{*}\omega$, where $\jmath:=\jmath_{\chi}\circ\jmath_{\scriptscriptstyle\Omega}:M_{\Omega}\rightarrow T^{*}Q$ defines the embedding of $M_{\scriptscriptstyle\Omega}$ into $T^{*}Q$.

### v2

@@ Abstract
The reduction of dynamical systems which are invariant under changes of global scale is well-understood, for classical theories of particles, and fields. The excision of the superfluous degree of freedom describing such a scale leads to a dynamically-equivalent theory, which is frictional in nature. In this article, we extend the formalism to physical models, of both particles and fields, described by singular Lagrangians. Our treatment of classical field theory is based on the manifestly covariant De-Donder Weyl formalism, in which the Lagrangian density is introduced as a bundle morphism on the pre-multisymplectic velocity phase space $J^{1}E$. The results obtained are subsequently applied to a number of physically-motivated examples, as well as a discussion presented on the implications of our work for classical General Relativity.
@@ I Introduction
Singular field theories are often amongst the most interesting class of models encountered in the description of natural phenomena [1, 2, 3]. Indeed, the most precisely tested mathematical theory to date, the Standard Model of Particle Physics, is a chiral quantum field theory, gauged under local transformations of the group $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$ [4, 5, 6]. Further, it is conjectured that any UV-complete theory of quantum gravity may possess no global symmetries, and that all symmetries must therefore be either spontaneously broken, or gauged [7, 8, 9]. It is thus clear that singular theories are of central importance to the advancement of modern theoretical physics.
Any attempt to rigorously develop a quantum theory requires a firm understanding of the underlying classical formalism. In this article, we present a framework in which singular theories that possess scaling symmetries may be reduced to a lower-dimensional, dynamically-equivalent description, that is frictional in nature [10]. For theories of particles, we utilise the pre-symplectic geometrical constraint algorithm developed in [11, 12, 13]. This is presented in section (III.1), following a review of elementary pre-symplectic geometry in section (II).
@@ II Geometrical Preliminaries
Suppose that $Q$ is an $n$-dimensional smooth manifold, corresponding to the configuration space of some mechanical system. A Lagrangian function $L:TQ\rightarrow\mathbb{R}$ is introduced on the tangent bundle, upon which local coordinates are denoted $(q^{i},v^{i})$ [35, 36, 37]. Throughout, the cases of greatest interest will be those in which the triple $(TQ,\omega_{L},E_{L})$ defines a pre-symplectic system [38]. In local coordinates, the closed 2-form $\omega_{L}$ and energy function $E_{L}$ are expressed as
$E_{L}=\frac{\partial L}{\partial v^{i}}v^{i}-L\quad\quad\quad\quad\omega_{L}=-\,d\theta_{L}=-\,d\left(\frac{\partial L}{\partial v^{i}}dq^{i}\right)$ (1)
@@
be of non-maximal rank, and thus singular. When studying the contact reduction of singular systems, we shall work almost exclusively within the Hamiltonian formalism. Our motivation for this is twofold: on the one hand, the Hamiltonian framework provides a well-defined bracket structure, which readily allows us to classify the constraint functions, and compute the dynamical evolution of any phase space variable. Further, not only does the Lagrangian formalism lack such a bracket, but the geometrical constraint algorithm introduces additional complications, which arise owing to a need to impose that, at each step, the solution satisfy the second-order problem. Since the dynamical content of the two formalisms is identical, we shall evade the undesirable features of the Lagrangian description by using it only as a means to obtain the corresponding Hamiltonian function.
@@
The dynamical problem is formulated via the introduction of a bundle morphism $\flat:T(T^{*}Q)\rightarrow T^{*}(T^{*}Q)$, with $\flat(X):=\iota_{X}\omega$; this is then restricted to $M_{0}$, with the resulting map being denoted $\flat_{0}$. We now seek a vector field $X_{H}$, such that
@@
Solutions to this equation generally do not exist on the entirety of $M_{0}$, thereby necessitating the implementation of a constraint algorithm. Such a procedure gives the maximal subspace of the cotangent bundle upon which the dynamical problem (2) possesses well-defined (albeit non-unique) solutions.
@@ III.1 The Hamiltonian Algorithm
Let $(M_{0},\omega_{0},H_{0})$ be the pre-symplectic Hamiltonian system, obtained from an almost regular Lagrangian $L$. It will often be convenient to describe $M_{0}$ as the zero-set of a collection of primary constraint functions $\{\phi^{a}\}$, with $a=1,\,\cdots,n-k$, in which $n=\textrm{dim}\,Q$, and $k=\textrm{rank}\,W_{ij}$ [40]. We shall assume that all subspaces generated are regular, closed submanifolds of $T^{*}Q$, whose natural injections are embeddings, and are denoted $\jmath_{i}:M_{i}\hookrightarrow M_{i-1}$. We therefore begin by seeking a vector field $X_{H}$ which satisfies
@@
Since $\omega_{0}$ is pre-symplectic, $\flat_{0}$ is not an isomorphism, and so we should only consider those points of $M_{0}$ at which $dH_{0}\in\textrm{Im}\,\flat_{0}$. From above, under the assumptions of reflexivity and topological closedness, we recall that $\flat_{0}(T_{x}M_{0})=T_{x}M_{0}^{\perp}$. However, $T_{x}M_{0}^{\perp}$ is precisely $\textrm{ker}\,(\omega_{0})_{x}$, and so the subset of $M_{0}$ of interest is
@@
We refer to these objects as the secondary constraint functions. In order for any vector field solution $X_{H}$, defined over the points of $M_{1}$, to be physically meaningful, it must remain tangent to $M_{1}$. If this were not the case, the solution would tend to evolve off the constraint surface, and thus cease to obey the restrictions placed upon the system. Tangency of the solution is not guaranteed, and so must be imposed as an additional constraint, leading us to consider only the following subset of points of $M_{1}$ [11]
@@
Vector field solutions on $M_{2}$ are now tangent to $M_{1}$, but, in general, not to $M_{2}$, requiring this condition to be imposed as a further restriction. It is then clear how the algorithm must proceed: at the $k^{\textrm{th}}$ step, the submanifold $M_{k}$ is composed of all points of $M_{k-1}$ at which the solution is tangent to $M_{k-1}$, but not necessarily to $M_{k}$, leading us to restrict $M_{k}$ to the smaller set $M_{k+1}$. This condition may be imposed requiring that $dH_{0}$ belong to the annihilator $(TM_{k}^{\perp})^{o}$, so that
@@
The constraint procedure may terminate in one of three ways [11]; however, the only dynamically-interesting eventuality is that there exists some $N\in\mathbb{N}$, such that $M_{N+1}=M_{N}$, with $\textrm{dim}\,M_{N}\neq 0$. Upon deducing the value of $N$ for which the algorithm terminates, we shall declare $M_{f}:=M_{N}$ to be the final constraint submanifold.
@@ III.2 Geometry of the Final Constraint Submanifold
Suppose that the constraint surface $M_{f}$, deduced from the algorithm above, is defined as the zero-set of the functions $\phi^{\alpha}$, with $\alpha=1,\,\cdots,D$. We introduce the matrix of Poisson brackets $J^{\alpha\beta}:=\{\phi^{\alpha},\phi^{\beta}\}$. If $\textrm{rank}\,J=D-F$, it follows that there exist $F$ $C^{\infty}$-linear combinations of the $\phi^{\alpha}$ which generate unobservable (gauge) transformations. [^1] Here, and throughout all subsequent analyses, we take the Dirac conjecture - that all first class constraints generate gauge transformations - to be true [41, 42, 43]. The goal is then to identify such combinations, corresponding to first class functions, and to modify the original set of constraints, so as to extract a maximal subfamily of second class functions. To this end, we introduce a ‘vector’ $v_{\alpha}=(v_{1},\,\cdots,v_{D})$, and demand that
@@
Having identified all such linearly independent combinations, indexing the resulting vectors via $v^{(a)}$, the $a^{\textrm{th}}$ first class constraint is simply $\Omega^{a}:=v^{(a)}_{\alpha}\phi^{\alpha}$. In order to extract the maximal subfamily of second class constraints, we seek DFD-F linearly independent vectors $w^{(j)}$, which are not in the kernel of $J$, so that
@@
Imposing the second class constraints $\chi^{j}=0$ as strong equalities, we obtain a hypersurface
@@
which inherits a closed 2-form $\omega_{\chi}$ which is symplectic. The Dirac bracket is precisely the algebraic structure required to ‘transfer’ this symplectic structure to functions defined on the whole of $T^{*}Q$. In particular, the dynamical evolution of any phase space function $f\in C^{\infty}(T^{*}Q)$ may be deduced from $\dot{f}=\{f,H_{T}\}_{D}$, in which $H_{T}:=H+\lambda_{a}\Omega^{a}$ is the total Hamiltonian; $\lambda^{a}$ are arbitrary Lagrange multipliers, and $H$ is any extension of $H_{0}$ to the full phase space.
@@
$M_{\scriptscriptstyle\Omega}$ is an embedded submanifold of $M_{\chi}$, whose 2-form $\omega_{\scriptscriptstyle\Omega}$ is maximally degenerate. Further, at each point $x\in M_{\scriptscriptstyle\Omega}$, the symplectic orthogonal $T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$ satisfies $T_{x}M_{\scriptscriptstyle\Omega}^{\perp}\subseteq T_{x}M_{\scriptscriptstyle\Omega}$, and so $M_{\scriptscriptstyle\Omega}$ constitutes a coisotropic submanifold of $M_{\chi}$ [45]. For such a class of space, the distribution $x\mapsto T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$ is involutive, and may locally be written as the span of the Hamiltonian vector fields $X_{a}$ corresponding to $\Omega^{a}$ via $\iota_{X_{a}}\omega_{\chi}=d\Omega^{a}$ [46]: that is to say
@@
Since the Hamiltonian vector fields of the first class constraints define an involutive distribution, we know that there exists a foliation of $M_{\scriptscriptstyle\Omega}$ into $F$-dimensional gauge leaves [47, 48]. The gauge orbits correspond to integral curves of the $X_{a}$, and they map points of $M_{\scriptscriptstyle\Omega}$ into other points, which produce dynamically indistinguishable configurations. Quotienting out by these gauge orbits, we obtain a physical phase space $\mathcal{P}$, which inherits a closed 2-form $\omega_{\scriptscriptstyle\mathcal{P}}$ that is symplectic.